consider the reaction A+B<----->C+D, which has an equilibrium constant K, equal to 3.4 x 10^2

1. If one begins a reaction by placing 0.6000 moles of A in 1.0 L container as well as 0.150 moles of B, what will be the equilibrium concentrations of A, B, C, and D?

2 Once the reaction in problem one reaches equilibrium, some additional B is injected into the flask from an outside source. LeChatelier's principal says the reaction will be

a) unchanged

b) shift to left

c) shift to right
Answer #1

0 Answer A + B (+ D Volum ( - IL n = 0.6000 mole ng = 0.150 mole [A] = c = 0.6000 = 0.6000 M [B] = 0.150 - 0.750 M лв с р toEquilibrium concentration = 0.6000 - 0.1499 LA leg LA Teg = 0.4501 M LB leg = 0.150 - 0.1999 = 0.0ool [Jeg = 0.1999 M [D7eg=